### v1

# Use as Many Surrogates as You Want: Selective Ensemble Attack to Unleash Transferability without Sacrificing Resource Efficiency

Abstract
In surrogate ensemble attacks, using more surrogate models yields higher transferability but lower resource efficiency. This practical trade-off between transferability and efficiency has largely limited existing attacks despite many pre-trained models are easily accessible online. In this paper, we argue that such a trade-off is caused by an unnecessary common assumption, i.e., all models should be identical across iterations. By lifting this assumption, we can use as many surrogates as we want to unleash transferability without sacrificing efficiency. Concretely, we propose Selective Ensemble Attack (SEA), which dynamically selects diverse models (from easily accessible pre-trained models) across iterations based on our new interpretation of decoupling within-iteration and cross-iteration model diversity. In this way, the number of within-iteration models is fixed for maintaining efficiency, while only cross-iteration model diversity is increased for higher transferability. Experiments on ImageNet demonstrate the superiority of SEA in various scenarios. For example, when dynamically selecting 4 from 20 accessible models, SEA yields 8.5% higher transferability than existing attacks under the same efficiency. The superiority of SEA also generalizes to real-world systems, such as commercial vision APIs and large vision-language models. Overall, SEA opens up the possibility of adaptively balancing transferability and efficiency according to specific resource requirements.
1 Introduction
Deep Neural Networks (DNNs) have been shown to be strikingly vulnerable to adversarial examples, i.e., maliciously crafted samples by adding small perturbations to benign inputs [35, 10]. An intriguing property of adversarial examples that makes them threatening in the real world is their transferability [1, 22, 45, 31, 16, 19, 47, 21], meaning that samples crafted on (known) surrogate models can also mislead (unknown) target models. Transferable attacks have been extensively studied, based on different strategies, such as gradient optimization, input transformation, and surrogate ensemble [49, 11].
Among them, surrogate ensemble is widely applied due to its simplicity and easy integration with other transfer techniques [40, 42, 50, 52]. In particular, there exists a well-known trade-off between attack transferability and the number of surrogate models, since using more models means lower resource efficiency regarding computational time and memory. Resource efficiency is even more critical when using transferable adversarial attacks for adversarial training-based defenses [38, 46, 34].
To avoid high resource overhead, the common practice is to simply restrict the number of models to only a few, typically $<5$ [6, 25, 30]. Under such a restricted condition, although complex, optimized ensemble strategies [44, 2, 37, 3] can somehow help, the resulting transferability is still far from using many more models, according to our experimental results in Section 4.2 and a concurrent study [27] on scaling law of transferability. In practice, it is meaningless to restrict the number of models since various pre-trained models are easily accessible online.
Figure 1: Our Selective Ensemble Attack (SEA) vs. conventional ensemble attacks (Ens) given $s$ easily accessible pre-trained models but restricted resources allowing only $m$ models per iteration. Our SEA($s$,$m$) dynamically selects $m$ diverse models across iterations compared to identical models in Ens($s$,$m$), leading to higher transferability. Here $m=3$ is used for illustration.
In this paper, rather than proposing a new ensemble strategy, we argue that the trade-off between transferability and resource efficiency can be removed, i.e., we can adopt as many surrogate models as we want to improve transferability without sacrificing resource efficiency. This is made possible by our new interpretation of model diversity: decoupling model diversity into within-iteration and cross-iteration dimensions. In general, both dimensions account for transferability, but only within-iteration diversity (corresponding to the number of models used in each iteration) accounts for resource efficiency. Specifically, we quantify these two dimensions of diversity based on gradient similarity [17, 20] to explore how they respectively contribute to the final transferability.
Based on the above new interpretation, we propose Selective Ensemble Attack (SEA), which dynamically selects a fixed number of diverse models (from easily accessible pre-trained models) across iterations. In this way, SEA increases cross-iteration diversity (by engaging more models across iterations) for higher transferability while maintaining within-iteration diversity (by fixing the number of models in each iteration). In contrast, the common practice, i.e., identical models should be adopted across iterations, blindly entangles the two dimensions of model diversity, and so it has to simultaneously increase transferability and resource costs. Figure 1 illustrates the difference between our SEA and conventional ensemble.
In sum, the main contributions of this work are as follows:
  • We argue that the trade-off between transferability and resource efficiency in model ensemble attacks can be removed. We attribute the problem to entangled within- and cross-iteration model diversity, caused by the common practice that forces identical models across iterations.
  • We propose Selective Ensemble Attack (SEA), which dynamically selects diverse models across iterations to leverage as many models as we want to unleash transferability without sacrificing resource efficiency.
  • Extensive experiments on ImageNet validate the superiority of our SEA in both transferability and resource efficiency over state-of-the-art ensemble attacks, including those rely on complex, optimized ensemble strategies.
2 Related Work
Various methods have been proposed to improve the transferability of adversarial examples, following similar principles in improving the generalizability of neural networks.
Model ensemble attacks. Model ensemble attacks generally combine outputs from multiple models. Specifically, Liu et al. [28] averages loss from multiple models.  Dong et al. [6] further proposes logits or predicted probabilities ensemble. Advanced ensemble strategies are also studied. Specifically, Xiong et al. [44] notices differences between ensemble models and proposes Stochastic Variance Reduction Ensemble (SVRE) attacks to stabilize the gradient update.  Chen et al. [2] proposes an adaptive ensemble attack method (AdaEA), which adapts the models according to their specific contributions.  Tang et al. [37] introduces a novel method of Stochastic Mini-batch black-box attack with Ensemble Reweighing (SMER) to boost the transferability.  Chen et al. [3] uncovers two common weaknesses of model ensembles and relies on them to design the common Weakness Attacks (CWA). Although model ensemble attacks notably boost adversarial transferability, their practical utility is constrained by the substantial computational resources and time required for simultaneous gradient calculations across multiple models.
Gradient optimization transfer techniques. MI [6] introduces momentum into adversarial attacks, stabilizing gradient update directions. NI [25] further integrates Nesterov accelerated gradient into transfer-based attacks to better handle the history and future gradients. VT [39] utilizes variance tuning to find a more stable direction for gradient updates. PGN [9] relied on penalizing gradient norm to achieve flat local maxima.
Input transformation transfer techniques. DI [43] applies random resizing and padding operations to the input during each attack iteration to increase sample diversity and mitigate overfitting. TI [7] optimizes adversarial perturbations using a set of translated images and also reduces computational costs by convolving gradients of untranslated images with a kernel matrix. SI [25] follows a scale-invariant strategy, which optimizes the adversarial perturbations over the scale copies of the input image. SSA [30] transforms input images into the frequency domain and uses multiple input copies with diverse frequency features.
In this work, following the common practice in model ensemble attacks [44, 3], we consider integrating the above gradient optimization and input transformation transfer techniques. Note that another type of transfer technique, model refinement [42, 49], is excluded since it focuses on model modification rather than model ensemble.
3 Methodology
3.1 Preliminaries of Model Ensemble Attacks
Given a classifier $f$ and benign input $x$ with ground-truth label $y$, $f(x)$ represents the logits output of classifier $f$ for input $x$. Naturally, the objective of adversarial attacks is to maximize $J(f({{x}^{adv}}),y)$ to craft an adversarial example ${x}^{adv}$ that is visually indistinguishable from $x$ but misleads the classifier to produce an incorrect output. Here, $J(f({{x}^{adv}}),y)$ is the loss function of the classifier, typically the cross-entropy loss function. Hence, adversarial example generation can be formalized as:
$\underset{{{x}^{adv}}}{\mathop{\arg\max}}\,J(f({{x}^{adv}}),y),s.t.||{{x}^{adv}}-x|{{|}_{\infty}}\leq\varepsilon,$ (1)
where $\varepsilon$ denotes the constraint on adversarial perturbations using the $L_{\infty}$ norm. Instead of directly solving the complex optimization problem of Equation 1, the Fast Gradient Sign Method (FGSM) [10] and its iterative version, I-FGSM [18], adopt the projected gradient descent.
The model ensemble is commonly used to improve adversarial transferability based on the assumption that adversarial directions learned from a diverse set of surrogate models can better generalize/transfer to unknown target models. Early studies on model ensemble attacks [28] propose to average the loss values of $J(f({{x}^{adv}}),y)$ in Equation 2, resulting in the ensemble loss function as:
$\underset{{{x}^{adv}}}{\mathop{\arg\max}}\,\sum\nolimits_{i=1}^{n}{\frac{1}{n}J({{f}_{i}}({{x}^{adv}}),y)},$ (2)
where $n$ denotes the number of ensemble models. Differently, logit output values can be averaged:
$\underset{{{x}^{adv}}}{\mathop{\arg\max}}\,\sum\nolimits_{i=1}^{n}{J(\frac{1}{n}{{f}_{i}}({{x}^{adv}}),y)}.$ (3)
In this paper, we adopt the logit ensemble variant due to its simplicity and effectiveness [6].
3.2 Decoupling Within-iteration and Cross-iteration Model Diversity
Model ensemble attacks are known to yield improved transferability but require significant time and memory consumption due to the gradient computations on multiple models. Consequently, existing work has commonly restricted the number of models (typically $<5$) for an acceptable resource efficiency [9, 2]. To fully investigate the impact of the number of models on the attack transferability, we conduct exploratory experiments by using 20 different models (see detailed descriptions in Section 4.1). As can be seen from the experimental results in Figure 2, using more models consistently improves the transferability, despite the performance saturation when too many models are adopted. This finding is supported by [15], which verifies that increasing the number of models in adversarial attacks can reduce the upper bound of generalization error in empirical risk minimization.
The above finding suggests that resource efficiency, which requires decreasing the number of models, may contradict transferability, which requires increasing the number of models. However, we point out that this contradiction is caused by the common practice that the models used in each iteration should be identical. To this end, we introduce the concept of cross-iteration and within-iteration model diversity. These two dimensions of model diversity have distinct effects on transferability and resource efficiency. On the one hand, both cross-iteration and within-iteration model diversity account for higher transferability. On the other hand, resource efficiency is only related to the number of models per iteration, corresponding to within-iteration model diversity. We present their formal definitions as follows.
Figure 2: Attack success rates (%) of MI transferring from an ensemble of $m$ models to eight different target models. The result for TI in the Appendix E shows similar patterns.
Definition 1 Cross-iteration and Within-iteration Model Diversity. Given an iterative attack, the model diversity achieved by the number of distinct models used throughout the entire process, $n$, is termed cross-iteration model diversity, denoted by $D_{c}$. The model diversity achieved by the number of distinct models used per iteration, $m$, is termed within-iteration model diversity, denoted by $D_{i}$. Here $m$ remains a constant in each iteration and $m\leq n$.
Quantifying diversity. To explore the relation between the two dimensions of diversity in Definition 1 and actual transferability, we further quantify the diversity. Inspired by existing work on transferable attacks/defenses, we quantify them based on gradient similarity [17, 20]. Intuitively, different models yield different gradient directions on the input image. In an ensemble attack, collective gradient directions from different models enable adversarial examples to better traverse the decision boundaries of multiple models, thereby improving the possibility of transferring them to unknown target models. For within-iteration diversity, we compute the cosine similarity $S$ between gradients from different models in the same iteration:
$D_{i}=\frac{2}{m(m-1)}\sum_{1\leq i<j\leq m}S(\nabla_{x}J(f_{i}(x)),\nabla_{x}J(f_{j}(x)))$ (4)
For cross-iteration diversity, we compute the cosine similarity between gradients from different models in two adjacent iterations:
$D_{c}=S(\sum\nolimits_{i=1}^{m}{\frac{\nabla_{x}J({{f}_{i}}({x_{t}}))}{m}},\sum\nolimits_{i=1}^{m}{\frac{\nabla_{x}J({{f}_{i}}({x_{t+1}}))}{m}})$ (5)
Table 1: Attack success rates (%) of MI when varying the number of cross-iteration models ($n$) and/or within-iteration models ($m$).
$n$ $m$ ATens4 ATens HGD RS Avg.
3 3 37.6 23.5 27.0 34.9 30.8
4 3 41.9 27.6 34.2 36.7 35.1
4 4 41.9 27.7 34.3 36.9 35.2
Table 2: Cross-iteration ($D_{c}$) and within-iteration ($D_{i}$) model diversity regarding gradient similarity for three attacks in Table 2.
$n$ $D_{c}$ $m$ $D_{i}$
grad sign(grad) grad sign(grad)
3 0.0023 0.0021 3 0.0062 0.0061
4 0.0022 0.0020 4 0.0055 0.0053
Table 2 shows the transferability results when varying the number of within-iteration models ($m$) and/or cross-iteration models ($n$). We can observe that solely increasing $n$ from 3 to 4 largely boosts the transferability. However, when further increasing $m$ from 3 to 4, the transferability is just improved a bit. This difference indicates the greater power of using as many models as possible across iterations (without sacrificing the resource efficiency). As expected, the diversity of the above three attacks quantified by gradient similarity in Table 2 is consistent with the results in Table 2.
3.3 Selective Ensemble Attack (SEA)
Algorithm 1 Selective Ensemble Attack (SEA) Algorithm
Input: A clean example $x$ with ground-truth label $y$; a classifier $f$ with loss function $J$; the logits of $m$ within-iteration models $l_{1},l_{2},...,l_{m}$; The set of available models across iterations $\mathcal{F}_{s}=\{f_{1},f_{2},...,f_{s}\}$.
Parameter: Perturbation size $\varepsilon$; iteration number $T$ and decay factor $\mu$.
Output: Adversarial example $x^{adv}$.
1:  $\alpha=\varepsilon/T$; $x=x_{0}^{adv}$; $g_{0}=0$;
2:  for $t=0$ to $T-1$ do
3:     Randomly select $m$ models from set $\mathcal{F}$:${\mathcal{F}}_{m}$ = $\{f_{1},f_{2},...,f_{m}\}$ = $random.sample(\mathcal{F}_{s},m)$;
4:     Input $x$ and output ${{l}_{i}}(x_{t}^{adv})$ for $i=1,2,...,m$;
5:     Fuse the logits as $l(x_{t}^{adv})=\sum\nolimits_{i=1}^{m}{{{l}_{i}}(x_{t}^{adv})}/m$;
6:     Get gradient as $g={{\nabla}_{x}}J(l(x_{t}^{adv}),y,\theta)$
7:     Update ${{g}_{t+1}}$ by ${{g}_{t+1}}=\mu\cdot{{g}_{t}}+\frac{g}{||g|{{|}_{1}}}$;
8:     Update $x_{t+1}^{adv}$ by $x_{t+1}^{adv}$ = $\text{Clip}_{x}^{\varepsilon}\{x_{t}^{adv}+\alpha\cdot\text{sign}({{g}_{t+1}})\}$;
9:  end for
10:  return ${{x}^{adv}}=x_{T}^{adv}$
Conventional model ensemble attacks with identical models per iteration enforce $n=m$ in Definition 1. As a result, even when there are many easily available models, they cannot use them to increase $D_{c}$ without increasing $D_{i}$ (i.e., requiring more resources). To address this limitation, we propose the Selective Ensemble Attack (SEA). SEA dynamically selects diverse models across iterations to leverage as many models as we want to increase $D_{c}$ while maintaining $D_{i}$ (i.e., maintaining the resource efficiency). SEA leads to overall higher diversity than the conventional ensemble methods, and as a result, improves the transferability.
Specifically, given $s$ easily accessible pre-trained models, denoted as $\mathcal{F}_{s}$ = {$f_{1}$, $f_{2}$, $f_{i}$, …, $f_{s}$}, SEA($s$,$m$) dynamically selects $m$ models in each iteration according to the required resource constraint. A larger $s$ generally leads to a larger number of cross-iteration models $n$, but $n$ may be smaller than $s$ when the iteration number is constrained (see detailed analyses of the expectation of $n$ in Appendix B). In contrast, the conventional method Ens($s$,$m$) has to limit $n$ to $m$. For model selection, we adopt the simple random selection and compare different sampling strategies in Table 9 of Section 4.3. The implementation details of SEA are presented in Algorithm 1.
4 Experiments
4.1 Experimental Setups
Dataset and models. The experiments are conducted on the ImageNet-compatible dataset [32], which is widely used for evaluating adversarial examples [6, 25]. This dataset consists of 1000 images from 1000 ImageNet categories (one image per category), which can be correctly classified by all the adopted models described as follows.
For the target (test) model, we consider 2 CNNs, 2 ViTs, and 4 Defenses. Specifically, 2 CNNs are Inceptionv4 [36] and Resnet-v2-152 [12]. 2 ViTs are ViT-base [8] and Swin-base [29]. 4 Defenses are ens4-adv-Inception-v3 (ATens4), ens-adv-Inception-ResNet-v2 (ATens) [38], HGD [24], and RS [5]. For the surrogate models, we select them from 20 pre-trained PyTorch models with 8 diverse architectures: resnet34/50/101/152 [12], vgg11/13/16/19 [33], densenet121/161/169/201 [14], fbresnet152 [12], dpn98/107/131 [4], nasnetamobile [53], senet154, and se_resnet101/152 [13]. Notably, the setting of 4 surrogate models is consistent with that of previous works [25, 39].
Figure 3: Attack success rate (%) of Ens(20,$m$) vs. SEA(20,$m$) when within-iteration model $m$ varies from 1 to 20. Results are reported for eight transfer baselines and averaged over eight target models.
Table 3: Attack success rates (%) of Ens vs. SEA with eight different transfer baselines. The upper bound using all 20 accessible models in each iteration is denoted as Ens(20,20), which costs ×$\times 5$ more computational time and memory than Ens(20,4)/SEA(20,4).
Baseline Attack CNNs ViTs Defenses Avg.
IncV4 Res152 ViT-B Swin-B ATens4 ATens HGD RS
Ens(20,20) 99.3 98.9 85.5 87.4 97.2 95.2 98.9 82.9 93.2
Ens(20,4) 89.7 91.2 54.7 52.0 78.3 69.1 86.4 59.6 72.6
MI SEA(20,4) 97.7 97.4 77.3 78.6 93.6 89.6 96.5 76.4 88.4
Ens(20,20) 99.4 98.9 84.7 86.3 96.5 93.5 99.1 83.6 92.8
Ens(20,4) 92.1 93.6 57.4 56.5 78.3 70.4 88.6 62.1 74.9
NI SEA(20,4) 98.9 97.8 73.7 76.1 91.8 87.9 97.7 73.1 87.1
Ens(20,20) 99.4 99.2 92.4 93.0 97.4 96.5 99.2 93.8 96.4
Ens(20,4) 95.3 95.1 75.0 73.6 89.2 85.4 94.6 79.3 85.9
VT SEA(20,4) 97.5 97.3 84.5 86.1 94.6 92.6 97.0 85.5 91.9
Ens(20,20) 98.9 99.2 95.6 93.2 98.5 97.9 99.3 98.2 97.6
Ens(20,4) 97.5 97.2 87.4 82.0 95.2 93.7 96.3 93.5 92.9
PGN SEA(20,4) 98.2 98.1 92.3 90.0 97.0 96.4 98.3 96.5 95.9
Ens(20,20) 99.4 99.4 92.9 91.6 98.1 97.6 99.3 95.9 96.8
Ens(20,4) 96.5 95.5 74.9 68.5 91.7 88.2 95.9 80.9 86.5
DI SEA(20,4) 98.0 97.3 80.9 77.2 95.2 93.0 97.0 85.8 90.6
Ens(20,20) 99.1 98.6 88.1 82.1 97.3 96.6 98.8 95.3 94.5
Ens(20,4) 90.5 91.3 64.9 51.0 84.9 79.6 89.7 80.0 79.0
TI SEA(20,4) 97.0 96.0 81.1 73.1 94.1 92.2 96.3 90.6 90.1
Ens(20,20) 99.6 98.9 84.1 86.2 96.1 94.0 99.2 83.5 92.7
Ens(20,4) 91.2 93.0 58.1 55.5 77.6 71.9 88.2 63.4 74.9
SI SEA(20,4) 98.4 98.2 75.0 75.8 93.2 87.8 97.5 72.9 87.4
Ens(20,20) 99.6 99.5 95.0 93.7 98.9 98.0 99.6 97.2 97.7
Ens(20,4) 98.6 98.2 85.2 81.2 95.3 94.5 97.1 90.6 92.6
SSA SEA(20,4) 99.1 98.8 92.1 90.4 97.5 96.4 99.1 93.8 95.9
Transfer baselines and hyper-parameter settings. We integrate conventional model ensemble attacks (Ens) or our SEA with eight popular transfer techniques that are based on gradient optimization (i.e., MI [6], NI [25], VT [39], and PGN [9]) or input transformation (i.e., DI [43], TI [7], SI [25], and SSA [30]). We also consider more complex ensemble strategies, such as SVRE [44], AdaEA [2], SMER [37], and CWA [3]. All the above transfer baselines have been introduced in Section 2.
For Ens and our SEA, the logit outputs of all models are averaged with equal weights. We set the perturbation budget $\varepsilon=16$, the number of iterations $T=10$. All integrated transfer techniques adopt the settings in their original papers. Specifically, for MI and NI, we set the decay factor $\mu=1.0$. For VT, we set the $\beta=1.5$ and the number of copies $N=20$. For PGN, we set the number of copies $N=20$, the balancing factor $\delta=0.5$, and the upper bound of neighborhood size $\zeta=3.0\times\varepsilon$. For DI, we set the probability $p=0.5$. For TI, we set the kernel length $k=7$. For SI, we set the number of copies $N=5$. For SSA, we set the tuning factor $\rho=0.5$ and the number of spectrum transformations $N=20$. For SVRE, AdaEA, SMER, and CWA, default settings are adopted.
Table 4: Transferability and resource efficiency of SEA vs. advanced ensemble strategies for (20,4).
Baseline Attack CNNs ViTs Defenses Avg. Time (min) Memory (GB)
IncV4 Res152 ViT-B Swin-B ATens4 ATens HGD RS
MI SVRE 94.5 94.5 64.7 62.5 84.7 78.0 93.1 66.3 79.8 36.8 15.5
AdaEA 91.0 91.3 56.6 53.5 78.4 70.6 87.7 61.8 73.8 88.7 19.1
SMER 95.3 95.4 65.3 61.0 85.6 79.2 92.6 68.8 80.4 92.9 21.9
CWA 94.9 94.8 55.8 47.9 75.9 65.7 87.7 63.3 73.2 9.1 20.1
SEA 97.7 97.4 77.3 78.6 93.6 89.6 96.5 76.4 88.4 5.9 14.1
TI SVRE 93.2 92.3 70.7 54.6 88.3 84.0 92.9 83.9 82.5 37.0 15.6
AdaEA 90.3 90.2 65.0 51.1 84.7 79.0 88.7 77.9 78.3 88.9 19.0
SMER 92.1 91.6 69.6 49.6 87.7 82.7 90.5 85.2 81.1 92.1 22.0
CWA 85.8 86.3 50.4 34.1 68.9 59.2 77.8 58.2 65.1 9.2 19.8
SEA 97.0 96.0 81.1 73.1 94.1 92.2 96.3 90.6 90.1 5.9 14.3
Evaluation metrics. We adopt attack success rate (%) to evaluate transferability and time consumption and memory usage to evaluate resource efficiency. The attack success rate measures the model’s misclassification rate on the 1000 adversarial examples from the dataset. Time consumption measures the time consumed to craft the 1000 adversarial examples from the dataset. Memory usage measures the maximum amount of GPU memory used during the attack. For attack success rate, higher is better, while for time consumption and memory usage, lower is better.
For clarity, we denote SEA and Ens under the ($s$,$m$) scenario as SEA($s$,$m$) and Ens($s$,$m$), respectively, where $s$ is the number of all easily accessible models and $m$ is the number of within-iteration models. When disregarding the architectural differences of models, the resource efficiency is directly proportional to $m$. For example, SEA(4,3) and Ens(4,3) are regarded as costing the same resource, and SEA(4,3) costs 1.5×1.5$\times$1.5 × resource of SEA(4,2). We also report the actual time and memory costs in Tables 4, 7, 8, 9 and in Appendix E, F, G. Additionally, the analysis of model loading time for SEA and Ens is reported in Appendix D.
Table 5: Attack success rates (%) of Ens vs. SEA on Google and Baidu Cloud Vision APIs with 100 randomly selected images.
Attack Google Baidu
Ens(4,3) 41 47
SEA(4,3) 44 51
Table 6: Attack success rates (%) of Ens vs. SEA on two commercial large vision-language models with 100 randomly selected images.
Attack Qwen ChatGLM
Ens(20,4) 40 39
SEA(20,4) 52 50
Figure 4: SEA adversarial images on Google and Baidu Cloud Vision APIs. The ground truth label of the original image is “snail”.
Table 7: Comparison of Ens vs. SEA on target attacks with DI-TI-MI transfer baseline.
Attack Surrogate models Target models Time (min) Memory (GB)
IncV3 IncV4 IncResV2 Res152 Avg. ATadv ATens3 ATens4 ATens Avg.
Ens(4,3) 86.4 86.3 78.6 79.2 82.6 13.0 12.9 34.3 28.3 22.1 38.5 18.3
SEA(4,3) 99.8 99.1 98.1 99.5 99.1 17.9 19.3 51.5 43.5 33.1 37.9 18.5
SEA(4,2) 100.0 99.0 98.2 99.5 99.2 16.9 17.9 47.3 39.5 30.4 25.7 16.1
Table 8: Transferability and resource efficiency of Ens vs. SEA with similar or diverse surrogates. In the similar cases, model ensemble is resnet18/34/50/101, vgg11/13/16/19, densenet121/161/169/201, or dpn68/98/107/131. In the diverse case, model ensemble is resnet101/vgg11/densenet161/dpn131. MI is the transfer baseline. Results for TI in the Appendix E show similar patterns.
Baseline Attack CNNs ViTs Defenses Avg. Time (min) Memory (GB)
V19 D121 ViT-B Swin-B ATens4 ATens HGD RS
Ens(4,4) 94.8 98.2 51.2 51.3 74.5 63.5 90.9 62.0 73.3 1.8 7.3
Ens(4,3) 91.8 97.6 46.4 42.6 68.6 55.3 84.7 58.1 68.1 1.5 6.9
resnet SEA(4,3) 95.1 98.1 51.3 49.2 74.6 63.0 90.5 61.3 72.9 1.5 6.9
Ens(4,4) 99.9 93.5 45.4 43.8 64.2 50.9 74.5 52.3 65.6 2.6 10.6
Ens(4,3) 99.9 90.9 40.9 38.6 57.7 43.9 67.2 50.4 61.2 2.0 10.1
vgg SEA(4,3) 99.9 93.4 44.6 42.2 63.3 48.9 74.8 52.1 64.9 2.1 9.8
Ens(4,4) 92.7 99.9 58.3 56.3 78.6 71.0 91.7 63.2 76.5 3.1 16.8
Ens(4,3) 91.3 99.8 52.1 51.3 73.3 66.1 87.8 60.6 72.8 2.4 13.6
densenet SEA(4,3) 93.3 100.0 57.5 57.4 78.9 71.8 90.3 63.7 76.6 2.4 13.8
Ens(4,4) 84.8 94.6 55.0 54.8 76.8 73.3 87.0 57.5 73.0 6.4 21.9
Ens(4,3) 81.8 93.5 50.5 50.5 73.1 67.4 83.5 55.2 69.4 5.2 20.3
dpn SEA(4,3) 85.1 95.0 53.1 56.2 78.6 72.9 87.9 58.2 73.4 5.0 20.2
Ens(4,4) 98.1 98.4 59.2 60.4 81.0 75.8 91.7 63.2 78.5 4.5 17.9
Ens(4,3) 94.9 97.6 51.9 53.1 75.6 67.8 86.1 58.5 73.2 3.5 16.0
diverse SEA(4,3) 97.9 98.3 58.0 61.0 80.1 74.7 91.3 63.7 78.1 3.5 16.1
4.2 SEA in Various Transfer Scenarios
Our SEA gives the potential to use many models without increasing the resource costs. To explore the potential of utilizing many models in SEA, we assume easy access to 20 pre-trained models from the PyTorch library. As can be seen from Figure 3, when varying the within-iteration model $m$ from 1 to 20, our SEA(20,$m$) consistently surpasses Ens(20,$m$) in all cases. The performance gap becomes gradually larger as $m$ decreases. This trend indicates that the superiority of our SEA stands out especially when the resource is restricted in practice. In particular, when the common $m=4$ is used, our SEA(20,4) outperforms Ens(20,4) by 8.5% on average. More detailed results for (20,4) on various target models are reported in Table 3, and results for (4,3), a more commonly used ensemble setting are in Table 12 of Appendix E.
We have the following observations. First, SEA performs better on all diverse (black-box) target models including CNNs, ViTs, and Defenses. This is because SEA involves more models across iterations (but without consuming more resources). Second, Ens(4,3) and SEA(4,3) consume very close resources, because they have the same within-iteration models. Third, Table 3 shows different transfer baselines consume different costs, and more costs generally lead to more effective attacks.
SEA vs. advanced ensemble strategies. We further compare our SEA to more complex, advanced model ensemble strategies, such as SVRE, AdaEA, SMER, and CWA. These strategies optimize the ensemble attack from different perspectives (see descriptions in Section 2), beyond equally averaging the logit or loss values of surrogate models, and follow Ens to use identical models across iterations. As can be seen from Table 4, our SEA still achieves the best transferability. For resource efficiency, the simplicity of the random model selection makes SEA consume substantially less computational time and memory. We also find that combining the above complex ensemble strategies with our SEA can further improve transferability. See detailed results in Table 14 of Appendix F.
Attacking real-world systems. We further compare Ens and SEA in attacking real-world systems: the commercial vision APIs, Google and Baidu Cloud Vision, and the large vision-language models (LVLM), Qwen and ChatGLM. Notably, given an (original or adversarial) image, we prompt the LVLMs using "Output a single-word label in English for this image." An adversarial image is considered successful when its output label differs from that of the original image. Table 6 and 6 show that SEA achieves better results than Ens on both APIs and LVLMs, with the MI transfer baseline. Figure 4 further shows successful SEA image examples on both APIs. Additional visualizations, also on two LVLMs, are provided in the Appendix I, where for LVLMs, we also try a finer-grained testing with the prompt "Describe the image no more than 50 words".
Targeted attacks. We further compare SEA and Ens under more challenging targeted attacks. Specifically, we follow the common practice [41, 48], using 100 iterations and a step size of 2, with DI-TI-MI as the baseline. The results in Table 7 demonstrate that SEA consistently outperforms Ens. For instance, SEA(4,2) achieves an 8.3% improvement in average attack success rate while reducing resource consumption compared to Ens(4,3), highlighting the superiority of our method.
4.3 Ablation Studies
Table 9: Attack success rates (%), Time consumption (min), and Memory usage (GB) of SEA under different sampling strategies. SEA(20,2) is used.
Base Sampling strategy ASR (%) Time (min) Memory (GB)
MI random w/ replacement 83.5 4.2 10.3
random w/o replacement 85.5 4.2 10.2
optimization 83.0 4.3 10.5
TI random w/ replacement 86.4 4.3 10.4
random w/o replacement 87.6 4.2 10.3
optimization 85.9 4.4 10.7
Figure 5: Attack success rate (%) of SEA ($s$, 4) by varying $s$ from 4 to 40.
Diversity of surrogate models. To validate the general effectiveness of our SEA, we consider ensemble settings with similar or diverse surrogate models. Specifically, we adopt four ensemble models from the resnet, vgg, densenet, or dpn family, as well as their combination, diverse. As can be seen from Table 8, our SEA(4,3) consistently achieves higher transferability than Ens(4,3) across different model combinations. In addition, using a more diverse combination of surrogate models yields generally higher transferability for both Ens and SEA. We also validate challenging cases with the ensemble of CNNs and ViTs in Appendix H, showing that even with such highly diverse model architectures, there are no opposite effects [2].
Model selection strategy. Selecting a subset of models from all models in each iteration results in an enormous search space. For example, when selecting 2 models from 20, there are $190^{10}$ possibilities over 10 iterations. Here we compare whether different selection strategies may have an impact on SEA. Specifically, we compare random (with or without replacement across iterations) selection and an optimization-based selection [51, 26], which updates the probability of each model being selected through differential augmentation. As can be seen from Table 9, random sampling without replacement yields the best results. This is because it avoids selecting the same models as those in previous iterations, leading to higher cross-iteration model diversity. For example, for SEA(20,2), random selection without replacement for 10 iterations yields the number of cross-iteration models $n=s=20$. The stability of random sampling is analyzed in Table 10 of Appendix C.
The number of accessible models $s$. Similar to Figure 3, we further conduct experiments on SEA(s,4) with varying $s$. Specifically, we fix $m=4$ and vary $s$ from 4 to 40 using the MI baseline to craft adversarial examples. As shown in Figure 5, the attack success rates steadily increase as $s$ increases. The growth is particularly notable when $s$ increases from 4 to 20. After $s$ exceeds 20, the improvement in attack success rates becomes slower, which is attributed to a certain level of saturation in the ensemble attack performance, as demonstrated in Figures 2 and 8.
5 Conclusion and Outlook
In this work, we point out the limitation of the existing model ensemble attacks in achieving a good trade-off between transferability and resource efficiency regarding time consumption and memory usage. To address this limitation, we propose to drop the common practice that models used per iteration should be identical. Concretely, we propose Selecting Ensemble Attack (SEA), which decouples the cross-iteration and within-iteration model diversity by dynamically selecting diverse models across iterations. In this way, SEA can use as many models as we want to improve transferability without sacrificing resource efficiency. Extensive experiments in various settings demonstrate the general superiority of our SEA. Overall, our SEA makes it possible to flexibly balance transferability and resource efficiency based on the actual resource constraint.
For future work, it is promising to further relax the constraint that the within-iteration models should be the same across iterations, possibly by dynamically selecting a different number of models per iteration. This will introduce more flexibility into our SEA, leading to a finer-grained adjustment of transferability vs. resource efficiency. It is also interesting to explore the generalizability of SEA to other domains beyond image classification.
Appendix A Limitations and Broader Impacts
In this work, we have explored and addressed the trade-off between transferability and resource efficiency. Another important property of (transfer-based) adversarial examples is the stealthiness of modified images, which has been simply restricted using an $L_{\infty}$ bound following existing work. In the future, we would explore whether it is possible to better trade off transferability and stealthiness.
Like other adversarial attack methods, if the proposed SEA method is maliciously used in practice, it may lead to security issues in deep learning models, representing one of its potential negative societal impacts. On the bright side, by leveraging the attack strength of adversarial examples, SEA can serve as a useful tool for evaluating and selecting the most robust models in safety-critical applications. In particular, the efficiency of our SEA would help speed up the evaluation and testing process. In addition, our study can provide researchers and practitioners constructive guidance for developing adversarial defenses.
Appendix B Analysis of the Expectation of $n$
We further quantitatively analyze the expectation of distinct models across iterations ($n$) used by Ens and SEA under the same conditions. Specifically, given the number of all easily accessible models $s$, the number of within-iteration models $m$, and the total iterations $T$, the expectation of distinct models across iterations is given by: $s\cdot\left(1-\left(\frac{s-m}{s}\right)^{T}\right)$. The derivation is as follows: For any given model, the probability of not being selected in a single iteration is $\frac{s-m}{s}$. Therefore, the probability that it is never selected over $T$ iterations is $\left(\frac{s-m}{s}\right)^{T}$, and the probability that it is selected at least once is $\left(1-\left(\frac{s-m}{s}\right)^{T}\right)$. By the linearity of expectation, the total expected number of distinct models selected over all iterations is the sum of the selection probabilities for each model, which yields: $s\cdot\left(1-\left(\frac{s-m}{s}\right)^{T}\right)$. As shown in Figure 7 and Figure 7, the expectation of $n$ in SEA increases as $s$ or $m$ increases. In contrast, the expectation of $n$ for Ens is always limited to $m$ since the same models are reused in each iteration.
Figure 6: The expectation of $n$ of Ens vs. SEA by varying $s$ from 4 to 40. The number of within-iteration models $m$ is 4.
Figure 7: The expectation of $n$ of Ens vs. SEA by varying $m$ from 4 to 40. The number of accessible models $s$ is 40.
Appendix C Stability Testing of Random Sampling in SEA
Table 10: The results (avg±std) of repeating Ens and SEA 10 times with the MI baseline.
Attack Vgg19 Dense121 ViT-B Swin-B
Ens(4,3) 73.9(±0.66) 71.1(±0.82) 26.3(±0.53) 26.8(±0.98)
SEA(4,3) 78.9(±0.61) 77.4(±0.33) 30.8(±0.41) 32.6(±0.99)
SEA(4,2) 78.4(±1.01) 77.2(±0.78) 30.1(±0.70) 31.9(±0.81)
We conduct an in-depth analysis of the stability of random sampling. Specifically, we repeat the experiments of SEA and Ens 10 times using the MI baseline. The experimental results in Table 10 show that Ens(4,3) and SEA(4,3) exhibit comparable and low variability, indicating that the random sampling strategy offers good stability.
Appendix D Model Loading Time
In this work, we load all available models at the beginning before the iterative attack optimization. This is fair for Ens and SEA because, given $s$ accessible models, Ens would also try to optimize the set of used $m$ models. Specifically, Ens would run the experiments multiple times, each time with a different set of $m$ models, and select the best run based on the transferability (between surrogate models). Table 10 confirms that such pre-selection is valuable since the results do vary across different sets of models. In contrast, our SEA does not require such a pre-selection but dynamically varies the set of models used in each iteration. Compared to GPU memories consumed for gradient calculation, the 4 models in SEA(4,3) only occupy 0.96 GB of memory, representing 4.04% and 1.22% of the total memory capacity on 3090 (24 GB) and A800 (80 GB) GPUs, respectively. Similarly, 20 models occupy 20.2% and 6.1% of the total memory capacity on 3090 and A800, respectively.
Furthermore, the results in Table 11 confirm that even with the additional time incurred by dynamic model loading, SEA still outperforms Ens in terms of attack effectiveness and efficiency in almost all cases when fewer models are used per iteration. Specifically, we measure the model loading time of 20 models used in SEA (20,4), and the average time is 3.7s, with variations from 2.5s to 6.1s. For an ensemble of 4 models and 10 iterations, the additional time of SEA is $3.7\times 4\times(10-1)=133.2$ (s).
Table 11: The comparison between Ens(20,4) with SEA(20,2).
Baseline Attack ASR(%) Time(min) Memory(GB)
Ens(20,4) 72.7 5.4 18.4
MI SEA(20,2) 82.4 5.3 11.2
Ens(20,4) 74.9 5.4 14.7
NI SEA(20,2) 82.2 5.3 11.3
Ens(20,4) 85.9 113.7 14.3
VT SEA(20,2) 87.9 59.2 11.6
Ens(20,4) 92.8 193.9 23.3
PGN SEA(20,2) 94.0 100.7 19.7
Ens(20,4) 85.8 5.5 14.3
DI SEA(20,2) 85.2 5.3 11.4
Ens(20,4) 78.9 5.5 14.3
TI SEA(20,2) 84.3 5.4 11.5
Ens(20,4) 74.8 27.3 14.4
SI SEA(20,2) 85.4 16.1 11.4
Ens(20,4) 92.6 113.0 15.8
SSA SEA(20,2) 93.7 60.1 11.5
Appendix E Additional Experiments in Varied Settings
Table 12: Attack success rates (%), Time consumption (min), and Memory usage (GB) of Ens vs. SEA with nine different baselines.
Baseline Attack CNNs ViTs Defense Avg. Time (min) Memory (GB)
V19 D121 ViT-B Swin-B ATens4 ATens HGD RS
Ens(4,4) 78.9 77.8 29.9 33.9 41.9 27.7 34.3 36.9 45.2 5.4 18.4
Ens(4,3) 75.0 71.5 27.1 28.5 37.6 23.5 27.0 34.9 40.6 4.5 13.3
SEA(4,3) 79.1 78.3 29.9 33.8 41.9 27.6 34.2 36.7 45.2 4.6 13.3
MI SEA(4,2) 78.9 77.7 29.7 32.7 41.9 27.7 31.0 35.9 44.4 3.1 11.2
Ens(4,4) 84.1 81.1 32.4 34.7 41.0 26.6 29.9 37.9 46.0 5.4 14.7
Ens(4,3) 79.3 76.9 28.4 28.3 35.6 22.0 25.1 36.6 41.5 4.5 12.4
SEA(4,3) 85.0 82.1 31.7 33.3 39.8 25.2 27.3 37.7 45.3 4.6 13.4
NI SEA(4,2) 84.9 81.7 31.4 31.3 36.2 24.7 23.9 36.1 43.8 3.1 11.3
Ens(4,4) 88.4 88.9 49.6 54.9 67.9 56.0 61.7 51.0 64.8 113.7 14.3
Ens(4,3) 86.3 86.0 44.3 49.6 63.0 50.8 55.8 46.5 60.3 85.9 12.6
SEA(4,3) 87.8 88.1 48.0 52.1 65.3 54.7 59.7 49.7 63.2 85.3 13.3
VT SEA(4,2) 86.2 85.8 45.3 50.5 62.8 48.6 57.7 46.1 60.4 57.0 11.6
Ens(4,4) 95.6 95.0 73.6 73.3 88.1 82.7 84.1 79.8 84.0 193.9 23.3
Ens(4,3) 93.6 93.9 69.0 67.3 84.4 78.8 79.3 75.6 80.2 142.4 21.0
SEA(4,3) 94.6 94.5 72.4 73.5 87.1 82.0 82.1 78.4 83.1 144.7 21.5
PGN SEA(4,2) 94.0 94.1 68.7 70.3 84.5 79.3 78.7 76.4 80.8 98.5 19.7
Ens(4,4) 90.3 91.5 50.5 53.2 66.6 52.3 63.7 50.6 64.8 5.5 14.3
Ens(4,3) 88.0 89.1 44.5 42.9 61.0 42.0 52.5 47.4 58.4 4.1 12.0
SEA(4,3) 90.0 90.4 48.3 48.6 64.9 50.3 61.4 49.5 62.9 4.6 13.3
DI SEA(4,2) 88.4 88.4 45.1 42.6 60.0 44.0 55.4 45.7 58.7 3.1 11.4
Ens(4,4) 82.7 84.1 41.0 36.9 59.4 49.6 62.8 46.7 57.9 5.5 14.3
Ens(4,3) 77.9 79.3 37.4 28.9 59.1 48.2 54.3 43.7 53.6 4.4 12.5
SEA(4,3) 82.7 82.6 40.9 35.9 59.6 49.5 62.5 46.6 57.5 4.6 13.2
TI SEA(4,2) 82.1 82.5 40.8 35.8 58.2 49.4 62.0 46.5 57.2 3.2 11.5
Ens(4,4) 94.5 95.9 53.3 53.8 75.7 58.9 66.1 57.3 69.4 27.3 14.4
Ens(4,3) 91.8 93.3 49.1 47.4 69.3 52.3 58.2 53.8 64.4 20.6 12.6
SEA(4,3) 94.0 95.9 52.5 51.9 75.1 56.6 64.6 56.2 68.4 20.2 12.6
SI SEA(4,2) 94.1 95.8 50.8 50.3 71.6 55.6 59.9 54.6 66.6 13.9 11.4
Ens(4,4) 95.8 96.3 71.5 72.2 85.8 77.1 81.1 72.5 81.5 113.0 15.8
Ens(4,3) 94.8 94.9 65.7 65.4 81.4 72.6 76.3 68.7 77.5 86.0 13.6
SEA(4,3) 95.7 95.1 69.8 70.2 84.3 74.4 78.8 71.5 80.0 84.5 13.9
SSA SEA(4,2) 94.8 94.1 66.9 66.8 81.7 72.0 74.1 67.5 77.2 57.9 11.5
Ens(4,4) 91.9 89.2 41.2 42.5 52.3 34.5 38.1 37.0 53.3 7.2 22.8
Ens(4,3) 87.6 86.1 34.4 32.6 45.5 28.6 31.9 34.5 47.7 4.8 21.5
SEA(4,3) 88.5 87.2 38.1 38.0 47.7 33.2 34.1 37.0 50.5 5.0 21.6
ghost SEA(4,2) 84.6 84.1 34.1 31.5 42.7 26.0 25.0 33.9 45.2 3.6 20.3
Figure 8: Attack success rates (%) of transferring from an ensemble of $m$ models to eight different target models. TI is the transfer baseline.
Table 13: Transferability and resource efficiency of Ens vs. SEA with similar or diverse surrogates. TI is the transfer baseline.
Surrogates Attack CNNs ViTs Defense Avg. Time (min) Memory (GB)
V19 D121 ViT-B Swin-B ATens4 ATens HGD RS
Ens(4,4) 96.5 99.1 61.8 48.7 84.0 76.7 90.4 82.4 80.0 1.9 7.2
Ens(4,3) 95.4 98.2 57.6 40.1 78.2 70.9 85.9 78.6 75.6 1.6 6.8
SEA(4,3) 96.9 98.8 61.1 46.8 83.2 76.6 90.9 82.3 79.6 1.5 6.8
resnet SEA(4,2) 96.1 98.8 59.7 43.7 81.4 74.0 88.7 82.3 78.1 1.2 6.4
Ens(4,4) 99.9 95.6 52.9 43.5 70.8 62.5 78.7 72.5 72.1 2.7 10.6
Ens(4,3) 99.8 94.4 49.4 38.7 67.8 58.0 72.2 68.8 68.6 2.1 10.1
SEA(4,3) 99.9 95.8 53.0 42.4 70.4 60.7 76.5 71.8 71.3 2.1 8.9
vgg SEA(4,2) 99.9 94.3 52.3 40.6 69.2 58.7 75.0 71.7 70.2 1.1 7.5
Ens(4,4) 94.7 99.9 66.0 53.5 85.1 81.9 92.5 81.4 81.9 3.2 16.3
Ens(4,3) 92.6 99.9 62.6 47.8 81.8 75.3 89.8 77.8 78.5 2.5 13.7
SEA(4,3) 94.3 99.9 64.7 52.3 84.8 80.8 91.8 81.6 81.3 2.5 13.6
densenet SEA(4,2) 94.0 99.9 63.5 48.2 83.6 79.0 90.4 81.1 80.0 1.9 10.7
Ens(4,4) 85.7 95.5 64.0 52.7 85.3 85.2 90.9 74.7 79.3 6.6 22.0
Ens(4,3) 83.2 94.5 58.1 46.4 80.1 77.1 86.0 72.1 74.7 5.2 19.9
SEA(4,3) 85.8 95.4 63.5 51.6 85.8 83.6 90.2 74.9 78.9 5.1 20.1
dpn SEA(4,2) 85.0 95.1 61.4 51.3 84.1 81.6 88.4 74.8 77.7 3.7 15.4
Ens(4,4) 97.3 98.3 68.2 57.9 87.2 83.5 92.8 82.5 83.5 4.5 17.8
Ens(4,3) 94.8 98.1 60.8 49.1 81.6 77.8 90.2 78.3 78.8 3.6 16.5
SEA(4,3) 97.2 98.6 66.3 56.4 86.5 83.1 92.2 82.1 82.8 3.5 15.9
diverse SEA(4,2) 97.1 98.5 63.7 54.7 85.4 81.9 91.0 80.8 81.6 2.6 14.1
Table 14: Transferability and resource efficiency of Ens vs. SEA integrated with four advanced ensemble strategies.
Base Ensemble strategy Attack CNNs ViTs Defense Avg. Time (min) Memory (GB)
V19 D121 ViT-B Swin-B AT3ens4 ATens HGD RS
MI Ens(4,4) 91.3 91.2 41.3 47.9 55.5 37.5 43.8 42.6 56.4 46.4 15.2
Ens(4,3) 87.0 86.1 34.9 38.4 45.2 30.7 35.4 39.1 49.6 33.9 14.8
SEA(4,3) 89.3 88.7 38.8 44.1 51.1 33.3 41.8 42.1 53.7 34.6 14.4
SVRE SEA(4,2) 85.7 84.9 34.0 37.5 45.3 28.6 34.5 37.2 48.5 23.2 14.1
Ens(4,4) 79.0 79.1 30.8 33.4 42.7 29.1 35.6 36.8 45.8 53.2 19.2
Ens(4,3) 74.8 71.4 26.2 27.1 32.6 20.7 21.9 32.0 38.3 37.6 14.2
SEA(4,3) 78.4 76.8 31.2 32.4 37.0 24.9 27.4 35.7 43.0 33.9 14.1
AdaEA SEA(4,2) 75.7 74.0 28.7 31.9 32.4 21.0 20.0 34.3 39.8 24.5 9.5
Ens(4,4) 91.7 92.4 41.7 48.1 53.9 36.2 40.8 43.2 56.0 100.3 15.2
Ens(4,3) 89.2 86.9 37.0 37.9 44.6 28.5 31.7 38.7 49.3 67.1 13.9
SEA(4,3) 90.0 88.3 37.8 39.5 47.5 31.0 34.9 39.4 51.1 65.9 13.3
SMER SEA(4,2) 83.9 82.8 32.5 31.9 41.8 25.4 29.9 37.7 45.7 33.7 12.3
Ens(4,4) 92.9 91.1 33.5 33.3 32.0 19.2 27.1 43.0 46.5 11.5 14.5
Ens(4,3) 85.5 80.1 26.5 23.3 26.8 16.3 25.2 37.9 40.2 8.7 13.5
SEA(4,3) 91.2 88.3 33.3 32.5 33.3 20.0 27.6 43.1 46.2 8.7 13.3
CWA SEA(4,2) 88.0 86.5 29.8 28.2 30.5 18.6 27.5 41.3 43.8 6.2 11.5
TI Ens(4,4) 92.0 93.2 53.7 45.1 74.0 66.3 78.7 56.3 69.9 44.7 15.4
Ens(4,3) 89.3 90.2 48.1 39.4 67.2 58.6 71.2 51.3 64.4 34.2 14.9
SEA(4,3) 91.1 92.3 53.1 44.8 73.2 64.5 78.5 57.2 69.3 34.3 14.5
SVRE SEA(4,2) 88.3 91.0 47.4 40.2 65.5 55.9 71.3 51.7 63.9 23.4 14.3
Ens(4,4) 83.2 84.2 42.1 35.5 60.0 51.2 63.6 47.3 58.4 57.3 19.2
Ens(4,3) 78.1 79.4 37.1 28.4 52.1 43.2 52.7 43.6 51.8 36.5 14.2
SEA(4,3) 83.4 83.5 40.9 36.8 59.2 50.0 60.2 46.7 57.6 35.5 14.1
AdaEA SEA(4,2) 80.4 80.9 40.2 30.6 54.4 46.9 54.9 43.7 54.0 23.8 9.4
Ens(4,4) 93.1 93.9 54.7 48.2 76.7 67.6 78.9 58.8 71.5 111.4 15.2
Ens(4,3) 90.0 90.3 47.4 39.1 65.9 57.2 68.4 51.2 63.7 67.4 14.1
SEA(4,3) 91.4 92.5 49.4 42.9 68.8 60.9 71.3 53.9 66.4 67.4 14.0
SMER SEA(4,2) 86.9 87.7 43.0 35.0 59.9 49.7 62.8 47.6 59.1 34.2 12.3
Ens(4,4) 89.5 86.5 33.2 28.3 39.0 27.3 26.2 41.2 46.4 11.3 14.5
Ens(4,3) 83.0 73.5 26.7 22.2 29.7 19.2 17.4 34.7 38.3 8.7 13.5
SEA(4,3) 89.1 86.9 31.9 27.7 38.2 26.5 26.6 41.4 46.0 8.6 13.3
CWA SEA(4,2) 85.6 84.6 31.8 26.9 38.1 25.9 26.5 41.4 45.1 5.9 11.6
We provide additional experimental results to demonstrate the validity of the conclusions presented in the main text. These experiments cover varied number of surrogate models, surrogate model architectures, and ensemble strategies. For experiments that have been reported with the MI baseline in the main text, we additionally evaluate TI, a popular input transformation-based transferable attack. We also add another popular surrogate model refinement attack baseline [23, 42, 49], ghost [23]. Ghost randomly drops out neurons in the surrogate model.
The results of these experiments are illustrated in Figure 8, Table 12, and Table 13. As can be seen, all the conclusions drawn from these additional results are consistent with those in Figure 2, Table 3, and Table 8.
Appendix F SEA Integrated with Advanced Ensemble Strategies
In the main text, we have compared our SEA to advanced model ensemble strategies: SVRE, AdaEA, SMER, and CWA. Here we further show that SEA can integrate them for further improvement of transferability. As can be seen from Table 14, our SEA consistently outperforms Ens by a large margin. For example, our SEA boosts the transferability of the state-of-the-art method, CWA, by 6.0% on average.
(a) MI
(b) TI
Figure 9: Time consumption and memory usage of Ens vs. SEA when varying the batch size. D-value represents the difference between Ens and SEA. MI and TI are the transfer baselines.
Appendix G The Impact of Batchsize on Resource Efficiency
The setting of optimization batch size may have an impact on the actual consumption of resources. To investigate this impact, we conduct experiments with the upper bound Ens(4,4) and our SEA(4,3) under varied batch sizes. As can be seen from Figure 9, as expected, our SEA(4,3) consistently reduces the costs of Ens(4,4). Specifically, the improvement is more sensitive to the memory usage than the time consumption. In addition, as the batch size increases, the difference in time consumption between these two attacks becomes smaller, but the difference in memory usage becomes larger. This means that in practice, one should adjust the batch size based on the actual constraints in time consumption vs. memory usage to better leverage the benefit of our SEA.
Appendix H Ensemble of ViTs and CNNs
Table 15: Attack success rates (%) of Ens vs. SEA with MI under 3CNNs vs. 3CNNs+1CNN/ViT model settings.
Attack Surrogate Model MobileV2 Dpn131 DeiT-B PiT-B
Ens Res101/Dense161/Vgg11 96.1 94.7 58.8 66.3
SEA +Res50 97.2 97.5 63.7 69.9
SEA +Dense201 96.9 97.3 66.9 73.8
SEA +ViT-B 97.1 95.9 84.4 77.9
SEA +Cait-S 96.8 96.1 93.8 84.0
We further validate the effectiveness of SEA under a more challenging ensemble setting involving both CNNs and ViTs, as opposite effects may occur when highly diverse models are selected [2]. Specifically, Ens ensembles three CNNs, whereas SEA adds a model with a different architecture to the ensemble. The results in Table 15 show that even when the ensemble models have diverse architectures (CNN and ViT), SEA still outperforms Ens on all target models. More specifically, incorporating an additional CNN (ViT) into the ensemble enhances the attack performance against CNN (ViT) target models.
Appendix I Visualizations of Adversarial Images on APIs and VLMs
In Section 4.2, the quantitative results on APIs and LVLMs (presented in Tables 6 and 6) have demonstrated that SEA outperforms Ens. Here, we provide visualizations of the original images and the adversarial images crafted by SEA, as shown in Figures 10, 11, 12, 13, and 14. It can be observed that even ensembling small image classification models can effectively attack APIs and LVLMs, posing significant challenges in the development of robust vision APIs and LVLMs.
Figure 10: Original images and SEA adversarial images on Google Cloud Vision.
Figure 11: Original image and SEA adversarial image on Qwen with the prompt "Output a single-word label in English for this image".
Figure 12: Original image and SEA adversarial image on ChatGLM with the prompt "Output a single-word label in English for this image".
Figure 13: Original image and SEA adversarial image on Qwen with the prompt "Describe the image, no more than 50 words".
Figure 14: Original image and SEA adversarial image on ChatGLM with the prompt "Describe the image, no more than 50 words".